OpenStudy (anonymous):

Find the area for each circle with a diameter of 6.5m. A. 30 m^2 B. 31 m^2 C. 33.2 m^2 D. 35.2 m^2

OpenStudy (anonymous):

area of circle = pi * radius^2 diameter = 6.5 m -> radius = 6.5/2 = 3.25 m Area = pi * 3.25^2 = pi * 10.5625 = 33.2 m^2

OpenStudy (anonymous):

it' c